OpenStudy (anonymous):

11x + 3y = 20 3x + y = 4 Using the two equations above, solve for x.

OpenStudy (anonymous):

its either x = 7 x = 5 x = 4 x = 9

OpenStudy (anonymous):

we-write 2nd equation as y=4-3x. plug in thi into the 1st equation. u get the x value.

OpenStudy (anonymous):

i have no idea how to do the problem can u just give me the answer please (:

OpenStudy (anonymous):

i better give u the way how to do it first. so u learn how to do problem like this. :)

OpenStudy (anonymous):

okay, but idk what you mean by plug it in to the equation

OpenStudy (anonymous):

idk how to do problems with two unknowns

OpenStudy (anonymous):

u have y=4-3x. so u substitute the y in the 1st equation as 4-3x st: 11x + 3y = 20 ---> 11x + 3(4-3x) = 20 solve the new equation to get x

OpenStudy (anonymous):

if u need to find both unknown. like in this case, after u get the x, substitute the value of x into any of the equation, then solve it to get the value of y.

OpenStudy (anonymous):

idk im just gonna guess i have no idea how to do it

OpenStudy (anonymous):

it was 4 (:

OpenStudy (anonymous):

i got it right

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yes. x=4.

OpenStudy (anonymous):

u r welcome.

OpenStudy (anonymous):

You must understand what these equations represent. Suppose you are given two equations in this form and are asked to solve for x and y: 11x + 3y = 20 ---------------- equation 1 3x + y = 4------------------------equation 2 Equations 1 and 2 represent two straight lines on a graph which has an x axis and a y axis. You are trying to find the point of intersection of these two lines on the graph. You can try it out using a graph paper and you will find the correct answer of the intersection point, namely (x,y)